# Derandomizing Multivariate Polynomial Factoring for Low Degree Factors

### v1

@@ 29      Let $z$ be a variable that $P$ depends on
30      Find the smallest $e\geq 1$ such that $P\not|\,\frac{\partial^{e}f}{\partial z^{e}}$ and add $(P,e)$ to list $L$
31return L
@@
The time complexity is now easy to see. Steps 2 and 3 take time $\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}$ and $\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$, respectively. The for-loop line 4 - 18 is executed $|\mathcal{H}_{{\boldsymbol{\alpha}}}|\leq\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$ times. Steps 5 - 17 take time ${\rm poly}(snd)$ by Lemma 2.4 and Theorem 2.6. Line 18 takes time ${\rm poly}(snd)\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$ to test irreducibility by Lemma 4.1, plus the time to check divisibility, $\textup{T}_{{\rm Div}(\mathcal{C}_{\mu}/\mathcal{C}_{\textup{sp}})}$.
In Line 20, since we have $P$ computed explicitely, we can choose a variable $z$ that occurs in $P$. Then we have at most $d$ divisibility tests with the derivatives of $P$ in Line 21. In summary, we get the time bound claimed in the theorem statement. ∎
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{{\mathcal{D}}})}=nd\,\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}+{\rm poly}(snd)\,\textup{T}_{{\rm Irred\mbox{-}Proj}({\mathcal{D}})}\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/{\mathcal{D}})}\,.$
@@ 4.1 Sparse factors of constant-depth circuits
A challenging open problem is to compute the sparse factors of a given sparse polynomial, or more generally of a given polynomial computed by a constant-depth circuit. Only exponential-time algorithms are known. It follows now that to make progress on this problem, it suffices to derandomize the irreducibility preserving projection of sparse polynomials in better than exponential time.
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{Depth-}{t}}|_{\mathcal{C}_{\textup{sp}}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}\,.$
@@
It follows that a subexponential bound on $\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$ would yield a subexponential algorithm to compute the sparse factors of a sparse polynomial, or, more general, of a polynomial computed by a constant-depth circuit.
Corollary 4.5.
If irreducibility preserving projection for sparse polynomials is in subexponential time, then the sparse factors of a sparse polynomial can be computed in subexponential time,
$\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}=(snd)^{O((sd)^{\epsilon})}\implies{\rm Factor}(\mathcal{C}_{\textup{sp}}|_{\mathcal{C}_{\textup{sp}}})=(snd)^{O((sd)^{\epsilon})}\,.$
@@ 4.2 Sum-of-univariate factors of sparse polynomials
Given a sparse polyomial, we show that its factors that are sums of univariates can be computed in polynomial time. The result was already shown by Volokovich [Vol15] with a different technique. We show that it also follows via Corollary 4.3.
We already know that PIT for sparse polynomials is in polynomial time. Moreover, Saha, Saptharishi, and Saxena [SSS13] showed that divisibility of a sparse polynomial by a sum of univariates is in polynomial time.
Theorem 4.6 ([SSS13, Theorem 5.2]).
@@
Theorem 4.7 ([SSS13, Theorem 5.2]).
@@
If $|\textup{var}(p)|\leq 2$, we can directly derandomize Theorem 2.11: Suppose $\textup{var}(p)=\{z_{1},z_{2}\}$ so that $p({\boldsymbol{z}})=p(z_{1},z_{2})$. There is a polynomial $P$ in $2\cdot 2=4$ variables of degree $2d^{5}$ such that any non-root of $P$ gives an irreducible projection of $p$, where we set the other variables of $p$ to zero, $z_{i}=0$ for $i=3,4,\dots,n$. Hence, the trivial hitting set (5) of size $(2d^{5}+1)^{4}=O(d^{20})$ can be used for PIT for $P$. Since we want to handle the blackbox case, we do not assume that we have $p$ in hand. Hence, we do not know whether actually $\textup{var}(p)=\{z_{1},z_{2}\}$. So finally we will try all $\binom{n}{2}$ possibilities to choose $2$ variables. For each possibility, say $\{z_{i},z_{j}\}$, we take the same hitting set, but put the values at $z_{i}$ and $z_{j}$ and set the remaining $z$-variables to 00. That way we get $\binom{n}{2}$ hitting sets, and take their union to get the final hitting set of size $O(n^{2}d^{20})$.
@@
Consider the substituting $\psi:z_{i}\mapsto\alpha_{i}x$, for $i\geq 4$. Then $\psi(p)\in\mathcal{C}_{\textup{SU}}(4,d)$, since $\sum_{i=4}^{n}p_{i}(\alpha_{i}x)$ is a univariate polynomial in $x$. Because $\textup{var}(\psi(p)|\geq 3$, polynomial $\psi(p)$ is irreducible by Theorem 4.7.
@@
$\widehat{p}(x,z_{1},z_{2},z_{3})=\phi(\psi(p))=p_{1}(\alpha_{1}x+z_{1})+p_{2}(\alpha_{2}x+z_{2})+p_{3}(\alpha_{3}x+z_{3})+\sum_{i=4}^{n}p_{i}(\alpha_{i}x)$
@@
Corollary 4.9 ([Vol15] Sum-of-univariate factors).
@@
5 Conclusion
We conclude with some open questions.
@@ 5 Conclusion
    Can we decide whether a given sparse polynomial is irreducible in deterministic subexponential time? The proof may already give a good bivariate projection that preserves irreducibility. Then Theorem 4.2 would give us a deterministic subexponential-time algorithm to find irreducible sparse factors of a sparse polynomial.
@@
    Can we find bounded individual degree sparse factors of a sparse polynomial (that is not restricted as bounded individual degree) in deterministic quasipolynomial time? Volkovich asked if multilinear factors of a sparse polynomial can be found in deterministic polynomial time [Vol15]. We do not even know a deterministic polynomial time algorithm to test if a sparse polynomial is divisible by a multilinear polynomial.
@@
    Can one compute all the factors of a sparse polynomial/constant depth circuit by constant depth circuits of small size? At least, can one find all the factors that are computable in constant depth? The recent result in [KRSV24] gives a deterministic subexponential-time algorithm that outputs a list of circuits (of unbounded depth and possibly with division gates) that includes all such factors, but there might be spurious factors as well in the list.
@@
    Given a blackbox computing the product of sparse irreducible polynomials $f_{i}$ with bounded individual degree, find $f_{i}$’s in deterministic polynomial time. [BSV20] gives a quasipolynomial time algorithm, when the input is sparse with constant individual degree and the factors are all sparse (polynomially upper bounded with respect to input polynomial’s sparsity).
@@ Acknowledgements
P.D. is supported by the project titled “Computational Hardness of Lattice Problems and Implications”, funded by National Research Foundation (NRF) Singapore. T.T’s research supported by DFG grant TH 472/5-2. A.S. was supported by DFG grant TH 472/5-1 when the work started.

### v2

@@ 30  Let $z$ be a variable that $P$ depends on
[…]
32return L
@@
The time complexity is now easy to see. Steps 2 and 3 take time $\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}$ and $\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$, respectively. The for-loop line 4 - 18 is executed $|\mathcal{H}_{{\boldsymbol{\alpha}}}|\leq\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$ times. Steps 5 - 17 take time ${\rm poly}(snd)$ by Lemma 2.4 and Theorem 2.6. Line 18 takes time ${\rm poly}(snd)+\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$ to test irreducibility by Lemma 4.1, plus the time to check divisibility, $\textup{T}_{{\rm Div}(\mathcal{C}_{\mu}/\mathcal{C}_{\textup{sp}})}$.
In Line 20, since we have $P$ computed explicitely, we can choose a variable $z$ that occurs in $P$. Then we have at most $d$ divisibility tests with the derivatives of $P$ in Line 21. In summary, we get the time bound claimed in the theorem statement.
@@
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\mu}|_{{\mathcal{D}}})}={\rm poly}(snd)\,(\textup{T}_{{\rm PIT}(\textup{Hom}[\mathcal{C}_{\mu}])}\,+\,\textup{T}_{{\rm Irred\mbox{-}Proj}({\mathcal{D}})}\,+\,\textup{T}_{{\rm Div}(\partial\mathcal{C}_{\mu}/{\mathcal{D}})})\,.$ (51)
@@ 4.1 Sparse factors of constant-depth circuits
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{Depth-}{t}}|_{\mathcal{C}_{\textup{sp}}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}\,.$ (52)
@@
After we submitted this paper, Bhattacharjee et al. [bhattacharjee2025closure] showed a subexponential bound on $\textup{T}_{{\rm Irred\mbox{-}Proj}(\mathcal{C}_{\textup{sp}})}$. It follows that the sparse factors of constant-depth circuits can be computed in subexponential time.
Corollary 4.5 ([bhattacharjee2025closure]).
$\textup{T}_{{\rm Factor}(\mathcal{C}_{\textup{Depth-}{t}}|_{\mathcal{C}_{\textup{sp}}})}=\left(n\,(sd)^{O(t)}\right)^{O((sd)^{\epsilon})}\,\,.$ (53)
In fact, Bhattacharjee et al. [bhattacharjee2025closure] showed that constant-depth circuits can be completely factored in subexponential time.
@@ 4.2 Sum-of-univariate factors of sparse polynomials
Given a sparse polyomial, we show that its factors that are sums of univariates can be computed in polynomial time. The result was already shown by Volokovich [volkovich2015deterministically] with a different technique. We show that it also follows via Corollary 4.3.
We already know that PIT for sparse polynomials is in polynomial time. Moreover, Saha, Saptharishi, and Saxena [saha2013case] showed that divisibility of a sparse polynomial by a sum of univariates is in polynomial time.
Theorem 4.6 ([saha2013case, Theorem 5.1]).
@@
Theorem 4.7 ([saha2013case, Theorem 5.2]).
@@
If $|\textup{var}(p)|\leq 2$, we can directly derandomize Theorem 2.11: Suppose $\textup{var}(p)=\{z_{1},z_{2}\}$ so that $p({\boldsymbol{z}})=p(z_{1},z_{2})$. There is a polynomial $P$ in $2\cdot 2=4$ variables of degree $2d^{5}$ such that any non-root of $P$ gives an irreducible projection of $p$, where we set the other variables of $p$ to zero, $z_{i}=0$ for $i=3,4,\dots,n$. Hence, the trivial hitting set (7) of size $(2d^{5}+1)^{4}=O(d^{20})$ can be used for PIT for $P$. Since we want to handle the blackbox case, we do not assume that we have $p$ in hand. Hence, we do not know whether actually $\textup{var}(p)=\{z_{1},z_{2}\}$. So finally we will try all $\binom{n}{2}$ possibilities to choose $2$ variables. For each possibility, say $\{z_{i},z_{j}\}$, we take the same hitting set, but put the values at $z_{i}$ and $z_{j}$ and set the remaining $z$-variables to 0. That way we get $\binom{n}{2}$ hitting sets, and take their union to get the final hitting set of size $O(n^{2}d^{20})$.
@@
Consider the substitution $\psi:z_{i}\mapsto\alpha_{i}x$, for $i\geq 4$. Then $\psi(p)\in\mathcal{C}_{\textup{SU}}(4,d)$, since $\sum_{i=4}^{n}p_{i}(\alpha_{i}x)$ is a univariate polynomial in $x$. Because $|\textup{var}(\psi(p)|\geq 3$, polynomial $\psi(p)$ is irreducible by Theorem 4.7.
@@
$\widehat{p}(x,z_{1},z_{2},z_{3})=\phi(\psi(p))=p_{1}(\alpha_{1}x+z_{1})+p_{2}(\alpha_{2}x+z_{2})+p_{3}(\alpha_{3}x+z_{3})+\sum_{i=4}^{n}p_{i}(\alpha_{i}x)$ (55)
@@
Corollary 4.9 ([volkovich2015deterministically] Sum-of-univariate factors).
@@
Conclusion
Finally, we collect some open questions.
@@ Conclusion
    Can we decide whether a given sparse polynomial is irreducible in deterministic quasi-polynomial time? The proof may already give a good bivariate projection that preserves irreducibility. Then Theorem 4.2 would give us a deterministic quasi-polynomial time algorithm to find the irreducible sparse factors of a sparse polynomial.
@@
    Can we find bounded individual degree sparse factors of a sparse polynomial (that may not have bounded individual degree) in deterministic quasi-polynomial time? Volkovich asked if multilinear factors of a sparse polynomial can be found in deterministic polynomial time [volkovich2015deterministically]. Unfortunately, we do not have any deterministic polynomial time algorithm even to test if a sparse polynomial is divisible by a multilinear polynomial.
@@
    Given a blackbox computing the product of sparse irreducible polynomials $f_{i}$ with bounded individual degree, find the factors $f_{i}$ in deterministic polynomial/quasi-polynomial time. Bhargava, Saraf, and Volkovich [bhargava2020deterministic] gave a quasi-polynomial time algorithm, when the input polynomial is sparse with bounded individual degree.
@@
    Finally, we highlight a question from [bhargava2020deterministic]. Given a blackbox computing the product of two (or more) unknown sparse irreducible polynomials, find those irreducible factors in deterministic quasi-poly or polynomial time.
@@ Acknowledgements
P.D. is supported by the SUG Grant (#025774-00001) Debordering and Derandomization in Algebraic Complexity funded by Nanyang Technological University. T.T’s research supported by DFG grant TH 472/5-2. A.S. was supported by DFG grant TH 472/5-1 when the work started.